A complex function can be modified by the equation: y = cos(x^3) 3x^2
Find the indefinite integral of the function ( ∫ cos(x^3)3x^2 dx) using a substitution method?
1
Expert's answer
2019-11-26T09:47:10-0500

Indefinite Integral of a function


We need to find the Indefinite Integration of a function using a Substitution method.


Solution:


Given,


"\int cos (x^3 ) \space 3x^2 \space dx"

"Let \space x^3 = t \space\\
 then \space d (x^3 ) = dt"

(since , d (xn ) = n . x (n - 1) dx)

"3 \times x^2 dx = dt"

Plug all these in the given Integral, then we get


"\int cos (x^3 ) \space 3x^2 \space dx = \int cos t \space dt"


"= \int cos t \space dt = sint + C"

re-plug the substitution,




"\int cos (x^3 ) \space 3x^2 \space dx = sin \space t + C = sin \space x^3 + C"

Answer:

"\int cos (x^3 ) \space 3x^2 \space dx = sin \space x^3 + C"